# LeapFrog: Accelerating Multiscale Materials Simulations with Machine Learning

### v2

@@ 1 Introduction
PF has been used to successfully tackle the formation of pure metals[1] and alloys[2], whether in traditional casting or welding processes [3] or the rapid solidification rates found in additive manufacturing processes [4], as well as in solid state precipitation reactions [5].
@@
Luckily, theoretical and software advances have been able to take advantage of the recent boom in high-performance computing. The Adaptive Mesh Refinement (AMR) algorithm developed by Greenwood et al. [6] dramatically improves the allocation of computational resources by dynamically increasing the system resolution only where it is needed to accurately evolve the solidification interface. Other efforts have been made to leverage hardware developments by integrating Graphical Processing Units into the simulation pipeline, to significant effect[7][8][9][10].
@@
Although these advancements have reduced the real-time costs of PF and expanded the system sizes and time scales it can access, the computational costs can nonetheless be prohibitive: most experimentally relevant simulations still require weeks to perform on modern computational clusters. For this reason, there is still the need for further exploration and development of tools that can continue to push PF modelling forward.
@@
Machine Learning is another tool that has gained attention, use, and development to great effect recently, and it has been demonstrated to be an apt complement to currently established computational tools [7, 6, 11, 12]. Specifically, in the field of materials design and solidification, Neural Networks (NN) have been successfully deployed in a generative capacity to simulations of amorphous carbon [13], but also to accelerate the evolution of systems undergoing spinodal decomposition[14][15], 1st order phase transitions within small toy systems[16], the directional epitaxial growth of polycrystals[17], and 3D grain evolution in additive manufacturing contexts[18].
@@
In this paper, we present a novel “LeapFrog” algorithm that combines AMR with a U-Net with a Convolutional Long-Short Term Memory (CLSTM) base trained with a multiscale loss metric that can accelerate the PF modelling of directional solidification of a dilute binary alloy. In Section 2 we outline the PF model used to generate the data, the design of the network’s architecture, and the specifics of the LeapFrog algorithm. In Section 3 we go through the quantitative checks on and adjustments made to the algorithm’s output and present the wall time savings achieved by the algorithm. Finally, in Section 4 we discuss very feasible further applications and expansions of the algorithm.
2 Methodology
@@
The section describes the details and phase field (PF) and neural network methodologies, the neural network (NN) algorithm, the NN training modality, and the way the phase field and NN algorithms are combined to form a unified simulation platform whose proof-of-concept is the purpose of this paper.
@@ 2.1 Phase Field Model
The Phase Field (PF) model generating the data that the neural network will train on and predict is that of a dilute binary alloy simulated through Model C with anisotropy and an enforced directional thermal gradient $G$ developed by Echebarria et al. (2004)[2] and pull velocity $V_{p}$. The reason for this choice of model is its capacity to quantitatively emulate the conditions and resulting solidification microstructure of many industrial processes such as casting. By exploring and tuning different combinations of $G\textrm{ \& }V_{p}$, specific patterns, length scales and materials properties can be selected. These microstructural patterns are result from playing out the following dynamical equations:
@@
The last line of Eq. 1 is the coupling of the phase field, moderated by the coupling constant $\lambda$ and the interpolation function $g^{\prime}(\phi\rightarrow\{0,1\})\rightarrow\{0,1\}$, to gradients in the concentration field $\Delta c$. The first term in the brackets in the third line is a measure of the deviation from the equilibrium chemical potential, the second term captures the temperature at a specific height $y$ along the thermal gradient $G$ imposed on the system at a specific time $t$ given the pull velocity $V_{p}$ and liquidus slope $m_{L}$. Eq. 2 is the time evolution equation of the concentration field where the first term on the right-hand-side modulates solute diffusion scaling by the liquid diffusion/mobility $D_{L}$, and interpolation function $Q(\phi\rightarrow\{0,1\})\rightarrow\{1,0\}$, and the gradient and Laplacian in the concentration and reduced chemical potential, respectively. The second term on the right-hand-side of Eq. 2 is the solute anti-trapping term introduced by Karma (2001)[19] that counteracts non-physical diffusion effects brought-on by the artificially wide interface width. Finally, the terms $\eta$ and $\nabla\cdot\vec{q}$ at the end of both dynamics equations are added non-conserved and conserved stochastic noise, used in field theoretic models to emulate the role of thermal fluctuations emergent at the atomic scale and which are washed out at the mesoscale.
@@
Simulated from Eq. (3) and initiated from the morphology in the left frame.
In this work, this above model is simulated on a 2048 by 1024 mesh with noise using the Adaptive Mesh Refinement algorithm developed by Greenwood et al. [6]. Multiple initial semicircular solid seeds are initialized at the top $y=0$ line and evolved 350 000 time steps, approximately 70 000 time steps past the point where the resulting dendrites reach the $y=2048$ mark, given the chosen pull velocity and time step resolution. These simulations required approximately 3.5 hours to run on 8 AMD Rome 7532 CPU cores at 2.4GHz with 256Mb L3 Caches.
Given the stochastic nature inherent in solidification microstructure evolution, as well as the fact that said microstructure does not display self-similar scaling in space and time, it would be overly ambitious to frame the task at hand as one of training a neural network algorithm with the above phase field model to predict the final late-stage microstructure directly from process parameters. The relation between process parameters and final microstructure is highly non-linear and would require an extremely large and comprehensive database for a neural network (NN) to learn from. Indeed, we feel that a more feasible, and practical, goal is to use machine learning to enable the use of larger, and adaptable, simulated time steps. Such an NN would complement traditional adaptive mesh algorithms, leading to dramatic reductions in computational complexity in both the space and time domains. As such, the problem posed to the neural network will be to effectively predict the following:
[…]
@@
As a proof of concept of this idea, this work will focus on the competitive dendritic growth regime of these systems and ignore the transient regime wherein the solidification front transitions from its initial condition to dendrite fingers. This regime sets the foundation for the complex final solidification microstructure that forms in practical alloys, as well as numerous other first order phase transformations driven by competitive cellular arrays. To make this criterion quantitative, phase field simulation data will over the time period ($t\in[65000,250000]dt$) will be used as samples for the training regiments of the machine learning algorithm proposed in this work.
2.2 Neural Network

### v3

# SupplementalMat_ML4MSPF.pdf

@@ 1 Introduction
PF has been used to successfully tackle the formation of pure metals[karmaRappelPF1996] and alloys[echebarriaQttvAlloy2004], whether in traditional casting or welding processes [dantzigsolidification] or the rapid solidification rates found in additive manufacturing processes [debroy2018additive], as well as in solid state precipitation reactions [zhu2004three].
@@
Luckily, theoretical and software advances have been able to take advantage of the recent boom in high-performance computing. The Adaptive Mesh Refinement (AMR) algorithm developed by Greenwood et al. [greenwood_quantitative_2018] dramatically improves the allocation of computational resources by dynamically increasing the system resolution only where it is needed to accurately evolve the solidification interface. Other efforts have been made to leverage hardware developments by integrating Graphical Processing Units into the simulation pipeline, to significant effect[yamanaka_gpu-accelerated_2011][guo_gpu-accelerated_2022][yang_gpu-accelerated_2017][kumar_ferrox_2023].
Although these advancements have reduced the real-time costs of PF and expanded the system sizes and time scales it can access, the computational costs can nonetheless be prohibitive. AMR aided PF is still not immune from the limitations explicit methods impose on the time steps that can be taken. Implicit solvers can bypass this limitation at small system sizes, but not at the system sizes required for experimentally relevant simulations due to the number of nodes required.
@@
As such, most experimentally relevant simulations still require weeks to perform on modern computational clusters. Moreover, multiple simulations at one set of process parameters is needed to obtain an accurate microstructure characterization of systems simulated with stochastic noise. For this reason, there is still the need for further exploration and development of tools that can continue to push PF modelling forward.
@@
Machine Learning is another tool that has gained attention, use, and development to great effect recently, and it has been demonstrated to be an apt complement to currently established computational tools [yamanaka_gpu-accelerated_2011, greenwood_quantitative_2018, sakane2022parallel, wang2022application]. Specifically, in the field of materials design and solidification, Neural Networks (NN) have been successfully deployed in a generative capacity to simulations of amorphous carbon [kilgour_generating_2020], but also to accelerate the evolution of systems undergoing spinodal decomposition[montes_de_oca_zapiain_accelerating_2021][hu_accelerating_2022], 1st order phase transitions within small toy systems[peivaste_machine-learning-based_2022], the directional epitaxial growth of polycrystals[qin_grainnn_2023], and 3D grain evolution in additive manufacturing contexts[choi_accelerating_2024].
While there are numerous other examples predicting the time evolution of second order transformations (e.g., spinodal decomposition), and limited examples of a first order transformations with a single field theory, there are presently few that consistently couple both an order parameters (interface motion) with diffusion-limited transport — essential to better predicting accurate final microstructure length-scales and alloy concentration profiles?.
There are significant differences, in terms of ML applications, between first and second order transformations. In 2nd order transformations, the universal scaling of space and time (e.g., domains growing according to $t^{\frac{1}{3}}$) present a clear pattern signal for a neural network to learn how spatial domains should evolve in time. No such universal scaling generally exists for first order transformations, and as a result there is a continuous need to maintain communication from the smallest scales (the interface width) to the higher length scales of thermal and solute diffusion that controls intercellular and side-branching spacing and morphology. Thus, exploration and development of ML practices that can address the multi-scale challenges of solidification problems are needed.
@@
In this paper, we present a “LeapFrog” algorithm that combines AMR with a U-Net with a Convolutional Long-Short Term Memory (CLSTM) base trained with a novel dynamic multiscale loss metric that can accelerate the PF modelling of the important case of directional solidification of a dilute binary alloy with thermal noise and two-sided diffusion [echebarriaQttvAlloy2004]. This algorithm is based on the original work of the thesis of the main author [Pinto_MSC]. In Section 2 we present the wall time savings achieved by the algorithm, as well as the checks that establish the quantitative fidelity of the algorithm’s output as well. In Section 3 we discuss very feasible further applications and expansions of the algorithm. Finally, in Section 4, we outline the PF model used to generate the data, the design of the network’s architecture, details of the multiscale loss, and the specifics of the LeapFrog algorithm.
2 Results
This section demonstrates the LeapFrog algorithm described in Section 4 and quantifies its time-acceleration features. We also review and quantify the fidelity with which the results of the LeapFrog algorithm match those of direct phase field simulations.
2.1 Acceleration
a) System Time vs. Wall Time comparison between a Phase Field (PF) simulation using Adaptive Mesh Refinement (AMR) (black) and the LeapFrog algorithm (composite) evolution of the same system in the dendritic regime.
b) Breakdown total Wall Time between PF+AMR, an “Idealized” scenario where the PF+AMR data is always organized in memory (minimizing lookup time), and the LeapFrog Algorithm (broken down by process).
c) Comparison of the resulting $\phi$ and d) $C$ system fields after being evolved with the PF+AMR and the LeapFrog algorithm introduced in this work.
A first example of the time savings of the LeapFrog algorithm are shown in Fig. 1. The data show the results of an application of LeapFrog with $N_{ML}=5\cdot 1000dt$ and $N_{PF}=2\cdot 1000dt$. It can be directly seen in panel a) that the overall wall time savings afforded by the LeapFrog algorithm increase the longer the system is evolved (in terms of system, i.e., physical, time). Even with the aforementioned overhead of the machine learning pipeline, the total evolution of the system, past some transient time, up to when the dendritic array reaches the end of the simulation domain, is achieved using less than half the wall time.
As noted in the latter portion of Section 4.2, this time comparison does not include the initial time needed to generate the training set and to train the initial network. During a parameter space exploration — the application for which the LeapFrog algorithm was designed for and where it could provide the largest benefit — multiple simulation runs per process parameter set must be generated to obtain an ensemble view of the solidification microstructure. As a result, the first simulations generated can double for both microstructure characterization and as a training set for the NN. Any subsequent simulation runs can then be accelerated using the trained network.
The pipeline just described is possible thanks to the network’s accessible 7-hour training time on a training dataset composed of 11 simulation runs (in addition to 3 for the testing set, and 1 for the validation).
It is noted that the efficiency of the LeapFrog algorithm is greater than an even “idealized” version of the Adaptive Mesh Refinement (AMR) algorithm (blue segments in panels a) & b)). By “idealized” we mean a version of the AMR that retains its parallelized asynchronous execution of calculations while spatially adjacent cells in the system are always assured to be adjacent, or at least close by, in memory.
In practice, results of computations are written to memory out of order in the standard AMR algorithm, which incurs an asymptotic increase in lookup time necessary to find a cell’s neighbours for gradient computation. An optimal time interval could be determined where the overhead needed to reset the locations of cells in memory would result in an overall speed-up. However, for our purposes, this is not done here for the standard AMR algorithm.
LeapFrog always benefits from a quasi-ideal performance of the AMR mesh because machine learning predictions are always written to AMR input files in order. This is seen in frame a) of Fig. 1, which shows that the slope of the AMR phase of the LeapFrog algorithm (data in blue) has a lower slope than the direct AMR algorithm, which as described above is not memory-cache optimized by default.
Average Wall Time per $1000dt$ of System Time
LeapFrog
MSCLSTM $2.24$\mathrm{s}$\$
MSCLSTM (With Overhead) $6.86$\mathrm{s}$\$
AMR
AMR $46.58$\mathrm{s}$\$
Ideal AMR $34.02$\mathrm{s}$\$
The overall microstructure evolution does present some differences between direct PF+AMR and LeapFrog-enabled PF algorithms. These mainly involve interface fluctuations affected by the presence of thermal noise. An example of this is seen in a dendrite extinction event between the last two dendrites on the right sides of frames c) & d) of Fig. 1. The formation of the halted (extinguished) tip appears slightly delayed between the two approaches, and the subsequent side-branch formation of the surviving dendrite arms is somewhat dampened in the LeapFrog algorithm. We expect that the discrepancy of the two approaches can be mitigated by balancing of $N_{ML}$ to $N_{PF}$ due to the absence of noise in the ML predictions. It is currently unclear how to incorporate to implement noise in the NN weights such as to, say, produce thermal fluctuations at interfaces, let alone satisfy the fluctuation-dissipation theorem.
2.2 Adaptive Time Stepping
To explore the adaptability of the LeapFrog algorithm in trading off between reducing simulation time and fidelity of produced solidification microstructure, another run is performed on the same initial seed as the one that produced the system in Fig. 1. This time, however, the algorithm is used with $N_{ML}=5\cdot 1000dt=N_{PF}$. The results are shown in Fig. 2.
a) System Time vs. Wall Time comparison between a Phase Field (PF) simulation using Adaptive Mesh Refinement (AMR) (black) and the LeapFrog algorithm (composite) evolution of the same system in the dendritic regime.
b) Breakdown total Wall Time between PF+AMR, an “Idealized” scenario where the PF+AMR data is always organized in memory (minimizing lookup time), and the LeapFrog Algorithm (broken down by process).
c) Comparison of the resulting $\phi$ and d) system fields after being evolved with the PF+AMR and the LeapFrog algorithm introduced in this work.
Going similarly through the panels, despite the less exploitative $N_{ML}/N_{PF}$ ratio, we can see in a) that we continue to obtain over a five-fold increase in simulation speed (per $1000dt$) leading to half the wall time needed to arrive at the end of the simulation domain. At the same time, we observe improved fidelity of side-branching and bulk/interface morphology well behind the solidification front, which demonstrates that the LeapFrog algorithm can be adjusted in its stepping ratio to adapt to the level of fidelity as required for the application.
We have found that the LeapFrog-simulated microstructure for the data of Fig. 1 remains true to the direct PF-simulated counterpart up to $t=301000dt$, which illustrates the network adequately performing well past the temporal bounds of its training set ($t=250000dt$). This suggests that the transformations that the network has learned to emulate PF-solidification in a generalized manner, not restricted to a location or time within the system’s evolution.
Additional details on the tradeoff between simulation accuracy and acceleration provided by the LeapFrog algorithm — as well as an additional application of this kind of tuning of the algorithm — can be found in Supplementary Note 3.
2.3 Prediction Quality of $\phi$ and $C$ Fields
The first numerical characteristics we examined are profiles of the concentration field (segregation). The main interest is verifying the concentration profiles: (i) along the core (i.e., center) of one of the developed dendrites arms, (ii) along isotherms within the liquid, and (iii) along isotherms within the solid. The bulk concentration values within the core of the dendrites were used to verify that the LeapFrog algorithm produces output consistent with known physics such as the Gibbs Thomson and flux boundary conditions at play at a moving solid-liquid interface, which are satisfied to an excellent degree by the present Phase Field (PF) model [echebarriaQttvAlloy2004]. The isotherm profiles, on the other hand, allow us to verify that the thermodynamics that underpin phase diagrams buried within the PF formulation is also respected.
Initially, simply taking the outputs of the LeapFrog (LF) algorithm, we obtain the profiles illustrated in Fig. 3. These exhibit small oscillations around the corresponding AMR solutions, as mentioned in Section 4.3.
We correct these oscillations by applying a smoothing filter to the area within the dendrites that are produced by the algorithm. This filter is implemented by producing a mask analogous to those used in feature weighted loss functions, illustrated in Fig. 4, and applying a unitary smoothing operator at the indicated areas.
b) The magnitude of the gradients in said concentration field.
This outputs concentration profiles almost exactly match the concentration profile predicted by direct PF simulations, as shown in Fig. 5.
Figure 5: Comparison of concentration profiles along a dendrite’s core (centerline) for a simulation evolved with Adaptive Mesh Refinement (AMR) and LeapFrog (LF) at $t^{*}=250000dt$. (red lines in Fig. 3) with smoothing mask illustrated in Fig. 4 applied to the $C$-field output from the LeapFrog algorithm. The core (centerline) concentration profile is shown in a). Also shown are the concentration profiles along the system’s isotherms in the liquid (b) and solid (c) phases.
Figure 6: Absolute of pixel-wise error between the AMR and LeapFrog ($5000dt$ in ML, $2000dt$ PF alternation) runs at $t=250000dt$ for a run that shows particularly increased side-branching.
Figure 7: Mean and standard deviation of $L^{2}$ Norm distance between the Power Spectral Density(PSD)/power spectra of the interface as defined by the order parameter and concentration fields, averaged over 10 different AMR and LeapFrog runs.
We also examine a pixel-wise difference for PF and LF runs for the case where there is a particularly high amount of side-branching is shown in Fig. 6. While the differences are small at any location other than side-branching sites, we do not expect them to ever be — or trend towards — zero at these regions. This is because the evolutionary history of dendritic growth with [physically consistent] thermal noise means that any two runs should never perfectly agree. Indeed, two frames in the pixel-wise comparison such as that in Fig. 6 do not necessarily indicate statistically invalid growth patterns; what is important is that both the PF and LF runs are part of the same time-evolution ensemble corresponding to a given set of process conditions (i.e., $G$ and $V_{p}$ in this case).
Another important way of quantitatively examining solidification microstructure is observing the power spectra of the interfacial profile and, in the case of directional solidification, the derived Primary Arm Spacing (PAS) [MikePRL, GurevichPRE]. Power spectra of the interface profile offer a better sense of the comparative distribution of the structures at different length-scales within the two systems. Given the stochastic nature of dendritic growth with thermal noise, comparing the power spectra between two “sibling” runs (i.e., runs that share the same initial 65 000 time steps, but that are then run one with PF+AMR and the other with LeapFrog) can be more informative than observing their pixel-wise differences.
To quantify the difference between PF+AMR and LeapFrog power spectra, or Power Spectral Densities (PSDs), and the latter’s comparative stability in predictions, we have calculated the $L^{2}$ norm distance between their respective PSD at each time step. Specifically, we average the mean & standard deviation over 10 pairs of “sibling” runs. The results are plotted in Figure 7. We can observe significant agreement between both run modalities since the average total $L^{2}$ norm distance between their PSDs remains, at the very most,  $10^{-2}$ — mostly staying within the  $10^{-3}$ order of magnitude.
There is a bulge around $t=150000dt$ in the left frame of Figure 7, which, upon further investigation, is caused by 2 out of the 10 runs in particular where dendrite extinctions occurred at different times in the PF and LF runs (see Supplementary Note 3 for further details). In fact, when these runs are removed from the sample, the $L^{2}$ norm curves seem to roughly plateau at the steady-state regime ($\sim 150000dt$, see Supplementary Note 3) and then increase sharply around the $200000dt$ mark.
The progressive increase observed in at late time in Fig. 7 corresponds to where the solidification front — and more importantly, the solute gradient preceding it — encounter the bottom boundary of the system. We hypothesize that the increase in the differences between the power spectra at this point can be attributed to the neural network encountering difficulties learning the specific growth dynamics of the system under finite size effects. These modified growth dynamics would only be observable in the final portions of the training runs, and occupy a comparatively smaller proportion of the training data than the steady state regime, which poses a greater challenge for the neural network to learn.
In Supplementary Note 3, we illustrate the cause of the first hump in the sensitivity of the $L^{2}$ norm between PSDs as reflecting statistical differences in microstructure that is otherwise consistent with the same process and material parameters. This metric remains a good approximate indicator of comparative stability, especially in its overall magnitude, but a further examination of its specific contributors are required to fully characterize its fidelity as a metric for assessing microstructure prediction fidelity.
As a further illustration of the usefulness of using PSD to characterize microstructure selection in directional solidification, we plot the average power spectra averaged over all sibling runs at their most divergent, i.e. $t=250000dt$, in Fig. 8 where we observe nearly perfect overlap. The main discrepancy here is at the shoulder around $k=025\cdot 8.33\times 10^{-6}\textrm{m}^{-1}$.
Observing the differences in Fig. 6 in conjunction with the analysis in Supplementary Note 3, this is most likely associated to the mismatch in dendrite extinction timing occurring around (x,y)(50,1125)(x,y)~(50,1125). It is assumed that they are a result of the different noise patterns brought on by the LeapFrog algorithm and, upon further examination (Supplementary Note 3), do not affect the solidification microstructure as far as its system morphology or length scale selection: There is nearly identical agreement in the ensemble predicted primary arm spacing.
It is noted that primary arm spacing (PAS) is specified through the $k$-mode of the highest peak in the power spectrum densities (PSD). As noted in Fig. 8, the PAS obtained from the PF+AMR and LF ensembles are in excellent agreement (an agreement that is present at any system time — see Supplementary Figure 15). These results further suggest that despite two instances of microstructure, at the same time, not being identical as derived from the PF+AMR-evolved system versus the LF system (or even from PF+AMR system alone), the statistical distribution from which system morphology is drawn appears to evolve in a stable and predictable manner. In other words, individual realizations evolve by exploring a statistical ensemble specified by a system process parameters $(G,V)$, system size and initial conditions.
Figure 8: Mean power spectra comparison with uncertainty intervals between direct Phase Field(PF)/AMR simulations (black) and LeapFrog(LF)-enabled PF simulation (red) for both the phase (left) and concentration (right) fields. Data is taken at $t=250000dt$. Red/Black barred areas designate regions of overlap between the PF and LF uncertainty. The power spectra have been constructed from averages over 10 simulations with different randomly initialized seeds, and the $k=0$ mode set to 0.
3 Discussion
We have presented a machine learning algorithm that is capable of accelerating multiscale phase field simulations of quantitative directional solidification of two-sided dilute binary alloys simulated with stochastic noise using Adaptive Mesh Refinement (AMR). This is achieved by training a neural network that combines Convolutional Neural Networks, Long-Short Term Memory networks and U-Net architectures to tackle the large system sizes containing numerous time-scales and a large range of length-scales inherent to the microstructure.
We have also developed a novel loss function for the training of the neural network that dynamically focuses learning on small-scale features of interest by developing cost-effective masks that highlight the locations where said features are present. These masks proportionally amplify small-scale feature contribution in the final loss metric. This approach is particularly beneficial because it allows for the targeting of very small or very faint patterns in the microstructure evolution, with minimal direct calculations based on the fields of the model. This becomes essential to any experimentally relevant pattern formation problems where small-scale interfaces are coupled to long-range diffusion to establish structure on larger scales.
We presented an adaptive time stepping algorithm that uses the above neural network (NN) architecture in tandem with a direct phase field (PF) simulator to generate microstructure predictions. These are combined in a “LeapFrog” (LF) fashion, whereby results from the NN are fed into the PF simulator — and vice versa. This LF algorithm is also versatile, since its utilization of ML and PF predictions can be tuned to either favor simulation acceleration or high fidelity with respect to stochastic and fine-grained features, depending on the needs of the application.
In its current prototypical form, where no explicit effort was made to reduce the overhead of passing information back and forth from the NN to the PF simulator, the combined simulation platform can provide a speed-up rate of at least 5–10 times faster than direct PF simulation. This speed is expected to become much larger when the overhead is factored out. The proposed pipeline could be beneficial when applied to process-microstructure characterization by significantly reducing the total computational time required for the exploration of parameter space — especially if practices such as transfer learning prove applicable.
The speed of the algorithm was also shown to not come at the cost of accuracy, as its predictions — insofar as microstructure statistics are concerned — exactly match those of traditional PF modelling. These include overall morphology, micro-segregation patterns, and primary spacing selection.
Various future directions and extensions of this work are possible. Most directly, the direct integration of the NN into the PF simulator would further increase computational speed up. Furthermore, training the network over a wider range of process conditions (e.g, $(G,V)$) could allow for interpolative predictions within regions of interest in phase space. This would provide the aforementioned speed-ups to simulation generation while bypassing the need for initial training data generation (i.e., at new $(G,V)$ points in our application).
Beyond that, the techniques presented in this work are also not explicitly tied to specific solidification paradigms. The dynamic loss function masking presented in Section  4.3 can be applied to aid the training of NNs on any system simulations that require interface tracking.
Finally, the network proposed in this work can also be applied to act as a local prediction tool in the so-called “mini-mesh” data structure at the heart of the finite difference scheme of the latest Adaptive Mesh Refinement (AMR) platform[greenwood_quantitative_2018]. This would allow a more direct integration of our NN into AMR, which would further enhance the experimentally relevant system sizes and times scales accessible to phase field modelling.
4 Methods
@@
The section briefly describes the details and phase field (PF) and neural network methodologies, the neural network (NN) algorithm, the NN training modality, and the way the phase field and NN algorithms are combined to form a unified simulation platform whose proof-of-concept is the purpose of this paper.
@@ 4.1 Phase Field Model
The PF model generating the data that the neural network will train on and then emulate is that of a dilute binary alloy simulated through Model C with anisotropy, two-sided diffusion, and an enforced directional thermal gradient $G$ developed by Echebarria et al. (2004)[echebarriaQttvAlloy2004] and pull velocity $V_{p}$. The reason for this choice of model is its capacity to quantitatively emulate the conditions and resulting solidification microstructure of many industrial processes such as casting. By exploring and tuning different combinations of $G\textrm{ \& }V_{p}$, specific patterns, length scales and materials properties can be selected. These microstructural patterns result from evolving the following dynamical equations:
@@
The last line of Eq. 1 is the coupling of the phase field, moderated by the coupling constant $\lambda$ and the interpolation function $g^{\prime}(\phi\rightarrow\{0,1\})\rightarrow\{0,1\}$, to gradients in the concentration field $\Delta c$. The first term in the brackets in the third line is a measure of the deviation from the equilibrium chemical potential, the second term captures the temperature at a specific height $y$ along the thermal gradient $G$ imposed on the system at a specific time $t$ given the pull velocity $V_{p}$ and liquidus slope $m_{L}$. Eq. 2 is the time evolution equation of the concentration field where the first term on the right-hand-side modulates solute diffusion scaling by the liquid diffusion/mobility $D_{L}$, and interpolation function $Q(\phi\rightarrow\{0,1\})\rightarrow\{1,0\}$, and the gradient and Laplacian in the concentration and reduced chemical potential, respectively. The second term on the right-hand-side of Eq. 2 is the solute anti-trapping term introduced by Karma (2001)[karmaPF2001] that counteracts non-physical diffusion effects brought-on by the artificially wide interface width. Finally, the terms $\eta$ and $\nabla\cdot\vec{q}$ at the end of both dynamics equations are added non-conserved and conserved stochastic noise, used in field theoretic models to emulate the role of thermal fluctuations emergent at the atomic scale and which are washed out at the mesoscale.
@@
Simulated by Eqs.(1) (3) and initiated from the morphology illustrated in $t=2000dt$.
In this work, this above model is simulated on the footprint of a 2048 by 1024 mesh with noise using the Adaptive Mesh Refinement algorithm developed by Greenwood et al. [greenwood_quantitative_2018]. Multiple initial semicircular solid seeds are initialized at the top $y=0$ line and evolved 350 000 time steps, approximately 70 000 time steps past the point where the resulting dendrites reach the $y=2048$ mark, given the chosen pull velocity and time step resolution. These simulations required approximately 3.5 hours to run on 8 AMD Rome 7532 CPU cores at 2.4GHz with 256Mb L3 Caches.
Given the stochastic nature inherent in solidification microstructure evolution, as well as the fact that said microstructure does not display self-similar scaling in space and time, it would be overly ambitious to frame the task at hand as one of training a neural network algorithm with the above phase field model to predict the final late-stage microstructure directly from process parameters. The relation between process parameters and final microstructure is highly non-linear and would require an extremely large and comprehensive database for a neural network (NN) to learn from.
In our opinion, a more feasible, and practical, goal is to use machine learning to enable the use of larger, and adaptable, simulated time steps. Such an NN would complement traditional adaptive mesh algorithms, leading to dramatic reductions in computational complexity in both the space and time domains. As such, the problem posed to the neural network will be to effectively predict the following:
$\Delta_{m}\phi(\vec{x},t_{n})\equiv\phi(\vec{x},t_{n+m})-\phi(\vec{x},t_{n})$ $=\int_{t_{n}}^{t_{n+m}}\left(\frac{\partial\phi}{\partial t}\right)dt$ (4)
$\Delta_{m}C(\vec{x},t_{n})\equiv C(\vec{x},t_{n+m})-C(\vec{x},t_{n})$ $=\int_{t_{n}}^{t_{n+m}}\left(\frac{\partial C}{\partial t}\right)dt$ (5)
@@
As a proof of concept of the above idea, this work focuses on the competitive cellular and dendritic growth regime of these systems and ignore the transient regime wherein the solidification front transitions from its initial condition to dendrite fingers. This regime sets the foundation for the complex final solidification microstructure that forms in practical alloys, as well as numerous other first order phase transformations driven by competitive cellular arrays. To make this criterion quantitative, phase field simulation data over the time period $t\in[65000,250000]dt$ is be used as samples for the training regiments of the machine learning algorithm proposed in this work.
4.2 Neural Network Architecture